OpenStudy (anonymous):

How Would I Solve The Systems " q-2r=4 " And " q+r=37 " ?

OpenStudy (amistre64):

which method do you prefer?

OpenStudy (anonymous):

Substution .

OpenStudy (amistre64):

i perfer elimination; easier on the eyes :)

OpenStudy (amistre64):

whats your sub then?

OpenStudy (anonymous):

Well We Can Do Elimination , Either One Is Fine , I'm Just Not Sure Which Or What Way Too Start .

OpenStudy (amistre64):

which is moot, they are both the same method just different paths :) lets do elimination so I can show you its finesse

OpenStudy (amistre64):

q - 2r = 4 q + r = 37 <-- notice that if we get this to be a +2r, it eliminates with the top

OpenStudy (amistre64):

q - 2r = 4 q + r = 37 <-- *2 q - 2r = 4 2q + 2r = 74 ------------- ; now we add em up 3q = 78 ; divide out the 3 q = 78/3

OpenStudy (amistre64):

simplify as wanted

OpenStudy (amistre64):

q - 2r = 4 q + r = 37 <-- notice now that when this is -q, the qs cancel

OpenStudy (amistre64):

q - 2r = 4 q + r = 37 <-- *-1 q - 2r = 4 -q - r = -37 ----------- ; add em up -3r = -33 ; divide off the -3 r = -33/-3 ; reduce as wanted

OpenStudy (anonymous):

Wait , Where Did They " 2q " Come From ?

OpenStudy (amistre64):

we use the properties of equations that tells us we can scale an equation to change its form, and it retains its same value

OpenStudy (amistre64):

(q + r = 37) 2 , we can multiply by 2 right? 2q + 2r = 74 is an equaivalent statement that has the same inherent value

OpenStudy (amistre64):

only now we can use it to our advantage

OpenStudy (anonymous):

Now I See , I Just Had Too Reread It .

OpenStudy (amistre64):

those answers reduce, so make sure you do that or else someone grading it might get upset :)

OpenStudy (amistre64):

you have any questions?

OpenStudy (anonymous):

So I Could Use Either 1 Or 2 To Multiply By And Still Get The Same Answer , Or No ?

OpenStudy (amistre64):

you can multiply by whatever number you want, but in the end you want something that is useful to you to use. sometimes you might need to mulitply the top and bottom equations by different numbers to get the to match up

OpenStudy (amistre64):

it takes practice to notice your easiest options; but in the end, you can brute math it dragging and kicking it all the way and still come out smelling like roses

OpenStudy (anonymous):

So I Stayed With The #2 And Multiplied , So I Got " Q=26" Which Equation Would I Use To Plug It Back In ? Too Check MySelf .

OpenStudy (amistre64):

either one; q = q = q = q so it doesnt matter :) id go with the q+r = 37 tho since its just simpler looking

OpenStudy (anonymous):

11 ?

OpenStudy (amistre64):

that is correct, good job!!

OpenStudy (amistre64):

im out of stickers but i got these: ♫♫♫

OpenStudy (anonymous):

Cool , Thanks But I Have Many More Problems That I Don't Know How To Do Or Really UnderStand , Can You Help Me With Them Too ?

OpenStudy (amistre64):

post them to the left so others can spot them and help out; i gotta be going to class soon

OpenStudy (anonymous):

Alright , Thanks .